OpenStudy (anonymous):

Using the given zero, find one other zero of f(x). Explain the process you used to find your solution. 1 - 2i is a zero of f(x) = x^4 - 2x^3 + 6x^2 - 2x + 5.

OpenStudy (anonymous):

once again we can find the polynomial with those zeros, pretty much in our heads now right?

OpenStudy (anonymous):

it is \[x^2-2x+5\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

your next job is to either divide or factor

OpenStudy (anonymous):

\[x^4 - 2x^3 + 6x^2 - 2x + 5=(x^2-2x+5)(quadratic)\]

OpenStudy (anonymous):

you can do it by division, or by thinking thinking is easier, because long division sucks

OpenStudy (anonymous):

plus, there is no good way to write long division here want to use the think method?

OpenStudy (anonymous):

So what exactly I am dividing is where Im at

OpenStudy (anonymous):

did you understand up to here \[x^4 - 2x^3 + 6x^2 - 2x + 5=(x^2-2x+5)(quadratic)\] ?

OpenStudy (anonymous):

Yes that (x^2 +1)(x^2 -2x +5) equals x^4 - 2x^3 + 6x^2 - 2x + 5

OpenStudy (anonymous):

oooh ok then you are done

OpenStudy (anonymous):

But then what is the zero ?

OpenStudy (anonymous):

actually the way i read the problem, it is actually much much easier than even that

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

"Using the given zero, find one other zero of f(x)"

OpenStudy (anonymous):

Would I plug that in for x?

OpenStudy (anonymous):

since \(1-2i\) is a zero, the other zero must be its conjugate \(1+2i\) done finished easy

OpenStudy (anonymous):

that is all it is asking for one zero is \(1-2i\) find another one all that work for nothing

OpenStudy (anonymous):

wooowwwwwwww if I would have thought of that a million years ago haha thanks again!!!!

OpenStudy (anonymous):

btw once you did all that work, you have all the zeros \(1+2i,1-2i\) and from \(x^2+1\) you get \(i\) and \(-i\) not that it asked ...

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

Oh okay thanks aton haha I gtg to bed now its 11:35 o.o !